# RedOne: Revealing Domain-specific LLM Post-Training in Social Networking Services

### v1

@@ Abstract
Fei Zhao, Chonggang Lu, Yue Wang, Zheyong Xie, Ziyan Liu, Haofu Qian, JianZhao Huang, Fangcheng Shi, Zijie Meng, Hongcheng Guo, Mingqian He, Xinze Lyu, Yiming Lu, Ziyang Xiang, Zheyu Ye, Chengqiang Lu, Zhe Xu, Yi Wu, Yao Hu, Yan Gao, Jun Fan, Xiaolong Jiang, Weiting Liu, Boyang Wang, Shaosheng Cao NLP Team, Xiaohongshu Inc., China caoshaosheng@xiaohongshu.com
@@ 4.1 Implementation details
[…]
@@ 4.3 Main Results
As shown in Tables 2 and 3, we conducted comparison between RedOne with baseline models across various tasks. Meanwhile, as illustrated in Figure 3, we compared RedOne with its base model across seven dimensions (six for general and one for SNS). Both results indicate that RedOne not only maintains strong general capabilities, even surpassing its base model on general tasks, but also exhibit exceptional effectiveness in the SNS domain. Additionally, RedOne achieves performance comparable to significantly larger models across most tasks, with limited improvement opportunities observed only in few areas. The results also demonstrate that scaling up RedOne consistently enhances performance over smaller versions, aligning with established model scaling laws. These findings underscore RedOne’s strong potential for further advances through continued increases in model size.
@@ 4.4 Task-specific SFT Comparison
To further explore the impact of base model selection on task-specific fine-tuning and validate our domain LLM’s effectiveness, we conducted experiments on two 7B-scale models: the original Qwen-2.5-Instruct (”Qwen”) and our SNS-adapted model (”RedOne”). We evaluated three variants: (1) Qwen-Finetuned, involving task-specific fine-tuning on Qwen; (2) RedOne-Finetuned, involving task-specific fine-tuning on RedOne; and (3) RedOne, representing zero-shot inference without further fine-tuning.
@@ 4.5 Ablation Study
In this section, we investigate the contributions of each stage in our training pipeline through comprehensive ablation studies, with results summarized in Table 5. The results demonstrate that introducing the CPT stage leads to a slight decrease in performance on general-domain tasks, while significantly boosting performance on SNS related tasks. This observation confirms that continue pretraining with domain-specific data effectively adapts the model to target domains, with the trade-off in general performance being minimal. Moreover, the sequential application of SFT and PO yields additional improvements, which indicates the effectiveness of our comprehensive training strategy.
@@ 4.6 Online Results
Harmful Content Detection Exposure Rate ($\downarrow$) $-11.23$
Post-View Search Click Page Rate ($\uparrow$) $+14.95$
@@ 5 Conclusion
Although our proposed method demonstrates high effectiveness, it requires extensive data processing, resulting in considerable resource costs. Additionally, while our approach shows promising results on existing SNS benchmarks, the limited availability of SNS-related benchmarks restricts our comprehensive understanding of the method’s performance across domain tasks.

### v2

@@ Abstract
Fei Zhao, Chonggang Lu, Yue Wang, Zheyong Xie, Ziyan Liu, Haofu Qian, JianZhao Huang, Fangcheng Shi, Zijie Meng, Hongcheng Guo, Mingqian He, Xinze Lyu, Yiming Lu, Ziyang Xiang, Zheyu Ye, Chengqiang Lu, Zhe Xu, Yi Wu, Yao Hu, Yan Gao, Jun Fan, Xiaolong Jiang, Weiting Liu, Boyang Wang, Shaosheng Caothanks: Corresponding author. NLP Team, Xiaohongshu Inc., China caoshaosheng@xiaohongshu.com
@@ 4.1 Implementation details
During the CPT stage, we follow the training process from Qwen2.5 (Yang et al., 2024a) over a mixed corpus of general and SNS-specific data. SFT is conducted for three epochs in step one and two epochs in step two, with a maximum sequence length of 16 38416\,384 using sequence packing, batch size of $128$, a linear warm-up ratio of 0.1. The learning rates are set according to model size: for the 7B model, we use $5\text{\times}{10}^{-6}$ in step one and $3\text{\times}{10}^{-6}$ in step two; for the 32B model, we use $3\text{\times}{10}^{-6}$ for both steps. Optimization is performed using AdamW Loshchilov and Hutter (2017) ($\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\epsilon{=}10^{-8}$). In the final PO stage, we employ a learning rate of $1\times 10^{-7}$, batch size of $64$, sequence length of $4096$, training for two epochs, with SFT loss coefficient set to 0.3.
@@ 4.3 Main Results
As shown in Tables 2 and 3, we conducted a comparison between RedOne with baseline models across various tasks in three categories. Meanwhile, as illustrated in Figure 3, we compared RedOne-7B and RedOne-32B with their base model across seven dimensions (six for general and one for SNS). Both results indicate that RedOne in all scales not only maintain robust general capabilities, even surpassing their base model on general tasks, but also exhibit exceptional effectiveness in the SNS domain. Additionally, RedOne achieves performance comparable to significantly larger models across most tasks, with limited improvement opportunities observed only in few areas. The results also demonstrate that scaling up RedOne consistently enhances performance over smaller variants, aligning with established model scaling laws. These findings underscore RedOne’s strong potential for further advances through continued increases in model size, as well as its promise for real world application.
@@ 4.4 Task-specific SFT Comparison
To further explore the impact of base model selection on task-specific fine-tuning and validate our domain LLM’s effectiveness, we conducted experiments on two 7B-scale models: the original Qwen-2.5-Instruct (“Qwen”) and our SNS-adapted model (“RedOne”). We evaluated three variants: (1) Qwen-Finetuned, involving task-specific fine-tuning on Qwen; (2) RedOne-Finetuned, involving task-specific fine-tuning on RedOne; and (3) RedOne, representing zero-shot inference without further fine-tuning.
@@ 4.5 Ablation Study
We also investigate the contributions of each stage in our training pipeline, with results summarized in Table 5. While CPT alone shows limited immediate gains, it establishes a crucial foundation for subsequent specialization. Based on the CPT model, adding SFT and PO brings average improvements of 0.55 and 1.90 on SNS-Bench compared to variants without CPT, indicating that CPT provides a stronger knowledge foundation for SFT to improve instruction-following and also offers a broader exploration space for PO. Additionally, both CPT and SFT lead to a performance drop on general benchmarks, whereas PO could effectively mitigate this decline and further enhance the overall results. Finally, the complete three-stage pipeline yields the strongest results on specialized benchmarks, with 66.88 on SNS and 48.11 on SNS-Trans, while maintaining competitive general-domain performance at 63.83, demonstrating the effectiveness of our training strategy.
@@ 4.6 Online Results
Harmful Content Detection Exposure Rate ($\downarrow$) 11.23-11.23
Post-View Search Click Page Rate ($\uparrow$) +14.95+14.95
@@ 5 Conclusion
Although our proposed method demonstrates strong effectiveness, it requires extensive data processing, resulting in considerable resource costs. Additionally, current models are still in a large scale, which increases online inference latency and serving expenses, limiting deployment in resource-constrained settings. Future work will explore lighter architectures, including model compression through quantization and distillation, as well as routing-efficient designs such as mixture-of-experts (MoE), to reduce latency and cost without sacrificing accuracy.
@@
Reproducibility
Due to data privacy and code security concerns, we are currently unable to fully release our datasets and code. However, our entire development pipeline is built upon widely adopted open-source projects and works Zheng et al. (2024); Rafailov et al. (2023); Dong et al. (2023), which ensures that users only need to organize their data according to the specified format in order to run it smoothly. In addition, we have reported detailed parameter settings used during training in Section 4.1, which further provides valuable references for the community to reproduce our pipeline.